Suppose that a good is in perfectly elastic supply at price of RM5.00. The market demand curve for this good is linear, with zero quantity demanded at a price of RM25.00. If the slope of this linear demand curve is -0.25 draw a supply and demand graph to illustrate the consumer surplus that occurs when the market is in equilibrium.
1
Expert's answer
2014-06-16T03:36:06-0400
If Ed = -0.25 and quantity demanded is zero at P = $25, then Qd = 25 - 0.25P.
As Qs = 5, in equilibrium Pe = $5. Qe = 25 - 0.25*5 = 23.75 units.
Consumer surplus = 0.5*23.75*(25 - 5) = $237.5.